# Semantic Iterative Reconstruction: One-Shot Universal Anomaly Detection

### 24 Mar 2026 v1

@@ Abstract
Unsupervised medical anomaly detection is severely limited by the scarcity of normal training samples. Existing methods typically train dedicated models for each dataset or disease, requiring hundreds of normal images per task and lacking cross-modality generalization.
We propose Semantic Iterative Reconstruction (SIR), a framework that enables a single universal model to detect anomalies across diverse medical domains using extremely few normal samples. SIR leverages a pretrained teacher encoder to extract multi-scale deep features and employs a compact up-then-down decoder with multi-loop iterative refinement to enforce robust normality priors in deep feature space. The framework adopts a one-shot universal design: a single model is trained by mixing exactly one normal sample from each of nine heterogeneous datasets, enabling effective anomaly detection on all corresponding test sets without task-specific retraining.
Extensive experiments on nine medical benchmarks demonstrate that SIR achieves state-of-the-art under all four settings — one-shot universal, full-shot universal, one-shot specialized, and full-shot specialized — consistently outperforming previous methods. SIR offers an efficient and scalable solution for multi-domain clinical anomaly detection. Code is available at https://github.com/jusufzn212427/sir4ad.
@@ 1 Introduction
To overcome these limitations, we propose SIR, a framework that trains a single universal model capable of detecting anomalies across diverse medical domains using few normal samples. The key innovation is to construct a small but highly heterogeneous training set by selecting exactly one normal sample from each of nine representative medical datasets, resulting in a total of only nine images. This per-dataset one-shot sampling strategy enables strong cross-modality generalization without requiring task-specific retraining.
SIR extends the reverse distillation paradigm by employing a pretrained teacher encoder to extract multi-scale deep features and a compact student decoder equipped with multiple iterative refinement loops within a single forward-backward pass. These loops progressively enforce robust normality priors on the deep representations. During inference, anomaly scores are computed through multi-scale cosine similarity aggregation followed by Gaussian smoothing and cross-loop fusion.
@@
  • We propose a hierarchical decoder that employs re-downsampling for semantic reconstruction on teacher features. This design enables more effective modeling of hierarchical normality in the deep feature space and delivers strong performance under few-shot and cross-domain settings.
  • We introduce a multi-loop iterative refinement mechanism for reverse distillation via semantic iterative reconstruction. This decoder-centric process significantly strengthens the normality constraint in deep feature space under the one-shot regime, yielding markedly improved robustness and cross-domain generalization.
  • Extensive experiments on nine diverse medical anomaly detection benchmarks demonstrate that SIR consistently outperforms previous few-shot and full-shot methods in image-level area under the ROC curve, establishing a new state-of-the-art under limited training data.
Figure 1: Overview of the SIR framework. The architecture consists of a frozen pre-trained teacher encoder $E$ and student decoder $D$. During training, the teacher encoder processes one normal image from each of multiple source domains, extracting multi-scale deep features. The student decoder then performs iterative semantic reconstruction on the most compressed features through recurrent loops. At inference time, for any previously unseen target domain, the same single model receives only a test image $x$, and the anomaly score and pixel-level anomaly map are finally obtained from the reconstruction discrepancies accumulated across all iterative loops.
@@ 3.2 Semantic Reconstruction Framework
This up-then-down architecture elevates conventional feature reconstruction into semantic reconstruction. The student decoder is optimized by minimizing the cosine-similarity reconstruction loss at both semantic levels:
@@ 3.3 Iterative Refinement in the Decoder
The core of SIR lies in its multi-loop iterative refinement mechanism, which works seamlessly with the semantic reconstruction process described in Section 3.2. Rather than performing only a single forward pass, the student decoder $D$ carries out $L$ iterative refinement loops inside each training iteration. The high-level feature $\boldsymbol{\phi}^{t}_{k-1}$ from the previous loop is fed back as input to the current loop (starting with $\boldsymbol{\phi}^{t}_{0}=\boldsymbol{\phi}^{t}$), creating a parameter-free recurrent refinement process.
@@ 4 Experiments
Table 1: AUC comparison on nine medical benchmarks.
Setting Method Source RSNA OCT2017 APTOS ISIC BraTS BR35H LAG VinCXR Brain Tumor
Full-shot specialized AE-U [29] MICCAI 86.5 - - 70.1 83.0 - 80.8 76.9 94.4
RD4AD [12] CVPR 74.7 97.3 90.2 68.3 74.8 99.6 67.5 57.8 93.2
EDC [18] TMI - 99.7 95.8 89.3 - 99.8 - - -
E2AD [35] TMI - 99.8 97.5 90.9 - 99.8 - - -
Ours - 99.6 99.9 100.0 95.2 99.2 100.0 100.0 97.3 99.9
Few-shot specialized D24FAD [13] LCLR 99.2 - 100.0 - - - 97.3 - 95.5
MVFA-AD [20] CVPR 84.9 99.3 89.6 - - - 81.0 - 96.3
INP-Former [28] CVPR 78.5 - 91.1 - - - 74.7 - 81.4
InCTRL [40] CVPR 82.7 - 89.5 - - - 71.1 - 91.8
AnomalyGPT [15] AAAI 62.7 - 77.6 - - - 57.5 - 78.9
Ours (One-shot) - 99.5 99.8 100.0 97.8 96.6 100.0 99.9 80.4 98.5
Universal (Ours) Full-shot universal - 99.6 99.8 100.0 94.0 95.0 99.9 99.9 74.4 97.4
Ten-shot universal - 99.6 99.5 100.0 85.4 92.8 100.0 99.9 85.3 99.2
One-shot universal - 99.2 97.9 99.8 84.1 71.6 91.6 98.2 67.9 96.5
@@
1 99.34 99.90 100.0 70.11 69.75 100.0 94.00 57.44 98.74 87.70
3 99.60 95.71 100.0 82.46 96.46 99.57 95.40 80.75 99.73 94.41
5 85.00 95.60 97.77 92.25 89.83 36.96 86.17 78.94 43.25 78.42
7 99.12 73.56 97.38 45.21 60.06 93.34 93.01 43.67 99.05 77.38
Table 3: Per-loop and fusion analysis ($L=3$).
@@
Loop 1 $f_{3}$ 97.98 95.24 99.72 30.60 86.63 88.39 96.91 40.44 92.96 80.99
Loop 1 $\phi$ 98.26 79.69 99.89 72.36 88.25 99.20 92.61 84.22 98.66 90.35
Loop 1 Fused 98.82 90.65 100.00 76.62 92.18 98.80 94.16 61.84 98.53 90.18
Loop 2 $f_{3}$ 98.67 92.98 98.93 55.57 92.28 43.02 85.12 63.41 96.75 80.75
Loop 2 $\phi$ 97.64 79.67 98.69 74.12 88.50 96.16 87.88 82.81 98.87 89.37
Loop 2 Fused 99.23 91.25 99.93 75.59 93.65 88.56 89.04 77.94 99.73 90.55
Loop 3 $f_{3}$ 92.21 87.01 98.66 53.59 88.91 31.47 87.94 77.26 92.35 78.82
Loop 3 $\phi$ 95.56 68.64 96.34 71.12 88.67 94.88 87.13 79.40 98.95 86.74
Loop 3 Fused 96.91 82.87 99.49 68.54 92.65 71.42 91.34 82.01 99.56 87.20
Final Fused Sum 99.60 95.71 100.00 82.46 96.46 99.57 95.40 80.75 99.73 94.41
Table 4: Best Performance with sufficient training data.
RSNA OCT2017 APTOS ISIC BraTS BR35H LAG VinCXR Brain Tumor
Best AUC 99.68 100.0 100.0 99.17 99.25 100.0 100.0 98.37 100.0
Figure 2: Qualitative visualization of anomaly maps generated by SIR under the one-shot universal detection setting. For normal samples, all four maps remain uniformly clean across loops. For abnormal samples, the anomaly regions become progressively sharper and more localized as the number of loops increases, with the final fused map providing the clearest boundary and highest contrast.
@@
RSNA VinCXR APTOS LAG
RSNA 99.52 52.19 - -
VinCXR 99.46 90.27 - -
APTOS - - 100.0 99.95
LAG - - 99.25 99.91
@@ 4.1 Experimental Setup
We evaluate SIR on nine publicly available medical anomaly detection benchmarks covering diverse imaging modalities and clinical tasks. The training and test set splits for all datasets follow the protocols used in prior works: RSNA [33] for pneumonia detection in chest X-rays follows AE-U [29]; OCT [23] for retinal diseases follows E2AD [35]; APTOS [4] for diabetic retinopathy grading follows E2AD [35]; ISIC [10] for skin lesion analysis follows MedIAnomaly [9]; Brain Tumor [2] for brain tumor detection in MRI scans follows MedIAnomaly [9]; BraTS [5] for brain tumor segmentation and classification in MRI scans follows MedIAnomaly [9]; BR35H [1] for brain tumor detection in MRI scans follows E2AD [35]; LAG [25] for glaucoma screening in fundus images follows MedIAnomaly [9]; and VinCXR [32] for lung anomalies in chest X-rays follows MedIAnomaly [9].
@@ 4.2 Main Results
Table 1 presents the image-level AUC (%) on the nine medical benchmarks. Under the most challenging one-shot universal setting, SIR achieves an average AUC of 89.7% across all datasets. This outperforms all five few-shot specialized baselines by substantial margins and even surpasses several full-shot specialized methods on multiple benchmarks. Notably, SIR maintains near-perfect performance ($\geq 99.5\%$) on OCT2017, APTOS, BR35H, and LAG despite using only nine normal images in total for training.
In the full-shot universal setting, SIR further improves its average AUC to 94.1%. In the ten-shot universal setting, the average reaches 93.3%. When evaluated in the conventional per-dataset full-shot specialized setting, SIR also delivers the highest scores on nine datasets and is very close to the upper bound achieved by dedicated full-shot models.
@@ 4.3 Effect of the number of iterative loops
We evaluate SIR with different refinement loop counts $L$. Table 2 shows the image-level AUC for each dataset as well as the average across all nine benchmarks. A single loop ($L=1$) yields 87.7% average AUC, indicating insufficient normality enforcement. Increasing to $L=3$ brings a substantial gain to 94.41%. Further increasing the number of loops leads to performance degradation ($L=5,7$). Thus, $L=3$ achieves the best accuracy-efficiency trade-off.
@@ 4.5 Best Performance with sufficient training data
For reference, when sufficient normal samples are available in the conventional per-dataset full-shot specialized setting, the best specialized models achieve the following image-level AUC on each dataset: RSNA 99.68, OCT2017 100.0, APTOS 100.0, ISIC 99.17, BraTS 99.25, BR35H 100.0, LAG 100.0, VinCXR 98.37, and Brain Tumor 100.0 (see Table 4). These values represent the practical upper bound achievable with abundant per-dataset training data.
@@ 4.6 Intra-domain cross-dataset one-shot generalization
To further examine generalization within the same imaging domain, we conduct targeted cross-dataset one-shot experiments on two modality pairs: chest X-rays (RSNA and VinCXR) and fundus images (APTOS and LAG). As shown in Table 5, when the single normal training sample is selected from the row dataset and the model is directly tested on the column dataset, SIR achieves near-perfect performance on the diagonal and strong transfer on the off-diagonal. These results confirm that SIR can effectively learn domain-specific normality priors from a single image and generalize robustly across different datasets within the same modality, demonstrating its practical potential for multi-institution clinical deployment without retraining.
These results demonstrate that the proposed semantic iterative reconstruction enables a single universal model to achieve strong cross-domain generalization and robust normality modeling under both extreme data scarcity and full-data conditions.
@@ 4.7 Anomaly Map Visualization
To qualitatively evaluate the localization capability of SIR, we visualize the anomaly maps generated under the one-shot universal setting. For each dataset, we randomly select representative normal and abnormal samples from the full test set. Following the exact inference pipeline, we compute four anomaly maps per sample: the fused map from Loop 1, Loop 2, Loop 3, and the final fused sum across all loops.
@@
Figure 2 presents typical visualization results. For normal samples, all four maps remain uniformly clean across loops, demonstrating that SIR correctly reconstructs normal semantics. For abnormal samples, the anomaly regions become progressively sharper and more localized as the number of loops increases, with the final fused map providing the clearest boundary and highest contrast. These visualizations confirm that the multi-loop iterative refinement not only improves quantitative AUC but also yields semantically meaningful and human-interpretable anomaly localization.
@@ 5 Conclusion
In this paper, we presented Semantic Iterative Reconstruction (SIR), a framework for unsupervised medical anomaly detection. SIR trains a single universal model on only one normal image from each of nine heterogeneous datasets, enabling direct anomaly detection across all corresponding test sets without task-specific retraining. The framework introduces a compact up-then-down decoder that performs semantic reconstruction of hierarchical teacher features, combined with a multi-loop iterative refinement mechanism that progressively strengthens normality constraints in deep feature space.
Extensive experiments across nine diverse medical benchmarks demonstrate that SIR achieves state-of-the-art image-level AUC under the extremely limited one-shot universal setting, consistently outperforming both full-shot and few-shot baselines. The method also maintains strong performance when more normal samples become available and delivers competitive results in conventional per-dataset scenarios. Ablation studies confirm that the iterative refinement and multi-scale fusion are essential for robust cross-domain generalization.
SIR offers a practical and scalable solution for real-world clinical applications where annotated data and large normal sets are unavailable.
@@
Although SIR achieves strong image-level detection through deep semantic reconstruction, its pixel-level localization remains limited by the inherent abstraction of feature-space modeling. Highly irregular lesion boundaries and subtle intensity variations may not be delineated with perfect precision. Future research could incorporate explicit pixel-level constraints, multi-scale attention mechanisms, or integration with foundation models to further refine localization accuracy while preserving the universal one-shot capability.

### 02 Apr 2026 v2

# Deep Semantic Iterative Reconstruction: One-Shot Universal Medical Anomaly Detection

@@ Abstract
Unsupervised anomaly detection in medical imaging is severely limited by the scarcity of normal samples and poor cross-modality generalization. Existing methods typically train dedicated models per dataset or modality, requiring hundreds of normal images and struggling to generalize across imaging protocols.
We present Deep Semantic Iterative Reconstruction (DSIR), a simple yet powerful convolutional distillation framework that trains a single universal model for robust anomaly detection across heterogeneous medical domains using extremely limited normal data. DSIR employs a frozen pretrained teacher encoder to extract multi-scale deep semantic features and a simple re-downsampling decoder that performs iterative semantic reconstruction via multiple refinement loops in a single forward pass. By mixing exactly one normal sample from each of nine diverse datasets, it learns strong cross-domain normality priors in deep feature space and generalizes effectively to unseen domains without task-specific retraining.
Extensive experiments on nine challenging benchmarks spanning chest X-rays, retinal fundus, skin lesions, and brain MRI show that DSIR achieves state-of-the-art performance across all four protocols—one-shot universal, full-shot universal, one-shot specialized, and full-shot specialized—with notable gains on difficult datasets.
@@ 1 Introduction
To overcome these limitations, we propose DSIR, a framework that trains a single universal model capable of detecting anomalies across diverse medical domains using few normal samples. The key innovation is to construct a small but highly heterogeneous training set by selecting exactly one normal sample from each of nine representative medical datasets, resulting in a total of only nine images. This per-dataset one-shot sampling strategy enables strong cross-modality generalization without requiring task-specific retraining.
DSIR extends the reverse distillation paradigm by employing a pretrained teacher encoder to extract multi-scale deep features and a simple student decoder equipped with multiple iterative refinement loops within a single forward-backward pass. These loops progressively enforce robust normality priors on the deep representations. During inference, anomaly scores are computed through multi-scale cosine similarity aggregation followed by Gaussian smoothing and cross-loop fusion.
@@
  • We propose a semantic reconstruction framework for one-shot universal medical anomaly detection.
  • We introduce a multi-loop iterative refinement mechanism that strengthens normality constraints in deep feature space, thereby enabling robust generalization and strong universality across diverse medical modalities with limited normal samples.
  • Extensive experiments on nine diverse medical benchmarks demonstrate that our method achieves state-of-the-art performance across one-shot universal, full-shot universal, one-shot specialized, and full-shot specialized settings.
Figure 1: Overview of the DSIR framework. The architecture consists of a frozen pre-trained teacher encoder $E$ and student decoder $D$. During training, the teacher encoder processes one normal image from each of multiple source domains, extracting multi-scale deep features. The student decoder then performs iterative semantic reconstruction on the most compressed features through recurrent loops. At inference time, for any previously unseen target domain, the same single model receives only a test image $x$, and the anomaly score and pixel-level anomaly map are finally obtained from the reconstruction discrepancies accumulated across all iterative loops.
@@ 3.2 Semantic Reconstruction Framework
This re-downsampling architecture elevates conventional feature reconstruction into semantic reconstruction. The student decoder is optimized by minimizing the cosine-similarity reconstruction loss at both semantic levels:
@@ 3.3 Iterative Refinement in the Decoder
The core of DSIR lies in its multi-loop iterative refinement mechanism, which works seamlessly with the semantic reconstruction process described in Section 3.2. Rather than performing only a single forward pass, the student decoder $D$ carries out $L$ iterative refinement loops inside each training iteration. The high-level feature $\boldsymbol{\phi}^{t}_{k-1}$ from the previous loop is fed back as input to the current loop (starting with $\boldsymbol{\phi}^{t}_{0}=\boldsymbol{\phi}^{t}$), creating a parameter-free recurrent refinement process.
@@ 4 Experiments
Table 1: Image-level AUC (%) comparison on nine medical anomaly detection benchmarks. “Specialized” denotes per-dataset training, while “Universal” indicates a single model trained on the mixed source domains. Best results are in bold.
Setting Method Venue RSNA OCT2017 APTOS ISIC BraTS BR35H LAG VinCXR Brain Tumor Avg.
Full-shot specialized RD4AD [12] CVPR 74.7 97.3 90.2 68.3 74.8 99.6 67.5 57.8 93.2 80.4
EDC [18] TMI 85.9 99.7 95.8 89.3 66.6 99.8 58.1 60.7 96.9 83.6
E2AD [35] TMI 87.5 99.8 97.5 90.9 63.6 99.8 67.2 59.4 97.6 84.8
SIR (Ours) - 99.6 99.9 100.0 94.6 99.0 100.0 100.0 97.7 99.9 98.9
One-shot specialized D24FAD [13] LCLR 99.2 - 100.0 - - - 97.3 - 95.5 -
INP-Former [28] CVPR 78.5 98.5 91.1 76.1 90.8 99.4 75.2 65.9 81.4 84.1
InCTRL [40] CVPR 82.7 91.6 89.5 81.6 89.0 - 80.0 60.8 91.8 83.4
SIR (Ours) - 99.6 99.9 100.0 98.8 97.2 100.0 99.9 91.4 98.4 98.3
Universal (Ours) Full-shot universal - 99.6 99.9 100.0 95.3 93.6 99.9 99.9 72.6 97.4 95.4
One-shot universal - 99.2 94.2 99.9 79.8 77.6 99.9 99.2 82.8 96.5 92.1
@@
1 99.37 100.00 99.99 82.58 64.76 96.59 80.32 73.95 97.45 88.33
2 99.58 96.58 100.00 73.22 60.50 98.09 79.68 90.76 98.79 88.58
3 98.90 99.02 99.99 93.02 69.76 100.00 99.08 94.56 95.82 94.46
4 99.57 96.34 99.64 83.18 85.84 100.00 99.76 94.71 97.68 95.19
5 99.63 96.62 100.00 72.67 53.26 100.00 98.62 49.00 97.50 85.26
Table 3: Per-loop and fusion analysis ($L=3$) – averaged over 3 random seeds (final iteration).
@@
Loop 1 $f_{3}$ 72.40 96.25 98.05 83.16 28.56 93.01 86.01 79.95 87.18 80.51
Loop 1 $\phi$ 98.41 90.23 91.26 74.33 66.91 99.26 95.09 85.28 94.35 88.35
Loop 1 Fused 96.84 94.54 99.99 91.46 51.67 97.70 94.26 85.12 94.12 89.52
Loop 2 $f_{3}$ 90.18 95.39 96.39 79.27 64.05 99.10 91.34 79.97 94.84 87.84
Loop 2 $\phi$ 97.50 94.62 86.96 76.63 69.49 98.86 95.67 87.49 94.59 89.09
Loop 2 Fused 97.02 96.77 99.64 83.01 69.54 99.70 95.96 87.11 95.56 91.59
Loop 3 $f_{3}$ 93.01 95.72 94.85 81.02 65.75 99.60 90.12 81.25 95.06 88.49
Loop 3 $\phi$ 96.16 94.98 84.36 78.61 69.62 98.70 95.47 88.36 94.51 88.97
Loop 3 Fused 96.44 97.30 98.62 83.40 70.59 99.66 96.02 89.32 95.45 91.87
Final Fused Sum 98.90 99.02 99.99 93.02 69.76 100.00 99.08 94.56 95.82 94.46
Figure 2: Qualitative visualization of anomaly maps generated by DSIR under the one-shot universal detection setting. For normal samples, all four maps remain uniformly clean across loops. For abnormal samples, the anomaly regions become progressively sharper and more localized as the number of loops increases, with the final fused map providing the clearest boundary and highest contrast.
@@
Train / Test RSNA VinCXR APTOS LAG BR35H Brain Tumor BraTS
RSNA 99.62 65.63 - - - - -
VinCXR 99.54 90.08 - - - - -
APTOS - - 100.0 99.94 - - -
LAG - - 89.96 99.87 - - -
BR35H - - - - 100.00 98.41 40.80
Brain Tumor - - - - 100.00 98.38 62.52
BraTS - - - - 97.58 75.07 96.73
@@ 4.1 Experimental Setup
We evaluate DSIR on nine publicly available medical anomaly detection benchmarks covering diverse imaging modalities and clinical tasks. The training and test set splits for all datasets follow the protocols used in prior works: RSNA [33] for pneumonia detection in chest X-rays follows AE-U [29]; OCT [23] for retinal diseases follows E2AD [35]; APTOS [4] for diabetic retinopathy grading follows E2AD [35]; ISIC [10] for skin lesion analysis follows MedIAnomaly [9]; Brain Tumor [2] for brain tumor detection in MRI scans follows MedIAnomaly [9]; BraTS [5] for brain tumor segmentation and classification in MRI scans follows MedIAnomaly [9]; BR35H [1] for brain tumor detection in MRI scans follows E2AD [35]; LAG [25] for glaucoma screening in fundus images follows MedIAnomaly [9]; and VinCXR [32] for lung anomalies in chest X-rays follows MedIAnomaly [9].
@@ 4.2 Main Results
Table 1 presents the image-level AUC (%) on the nine medical benchmarks. Under the most challenging one-shot universal setting, DSIR achieves an average AUC of 89.7% across all datasets. This outperforms all five few-shot specialized baselines by substantial margins and even surpasses several full-shot specialized methods on multiple benchmarks. Notably, DSIR maintains near-perfect performance ($\geq 99.5\%$) on OCT2017, APTOS, BR35H, and LAG despite using only nine normal images in total for training.
In the full-shot universal setting, DSIR further improves its average AUC to 94.1%. In the ten-shot universal setting, the average reaches 93.3%. When evaluated in the conventional per-dataset full-shot specialized setting, DSIR also delivers the highest scores on nine datasets and is very close to the upper bound achieved by dedicated full-shot models.
@@ 4.3 Effect of the number of iterative loops
We evaluate DSIR with different refinement loop counts $L$. Table 2 shows the image-level AUC for each dataset as well as the average across all nine benchmarks. A single loop ($L=1$) yields 87.7% average AUC, indicating insufficient normality enforcement. Increasing to $L=3$ brings a substantial gain to 94.41%. Further increasing the number of loops leads to performance degradation ($L=5,7$). Thus, $L=3$ achieves the best accuracy-efficiency trade-off.
@@ 4.5 Best Performance with sufficient training data
For reference, when sufficient normal samples are available in the conventional per-dataset full-shot specialized setting, the best specialized models achieve the following image-level AUC on each dataset: RSNA 99.68, OCT2017 100.0, APTOS 100.0, ISIC 99.17, BraTS 99.25, BR35H 100.0, LAG 100.0, VinCXR 98.37, and Brain Tumor 100.0 (see Table LABEL:tab:best_performance). These values represent the practical upper bound achievable with abundant per-dataset training data.
@@ 4.6 Intra-domain cross-dataset one-shot generalization
To further examine generalization within the same imaging domain, we conduct targeted cross-dataset one-shot experiments on two modality pairs: chest X-rays (RSNA and VinCXR) and fundus images (APTOS and LAG). As shown in Table 4, when the single normal training sample is selected from the row dataset and the model is directly tested on the column dataset, DSIR achieves near-perfect performance on the diagonal and strong transfer on the off-diagonal. These results confirm that DSIR can effectively learn domain-specific normality priors from a single image and generalize robustly across different datasets within the same modality, demonstrating its practical potential for multi-institution clinical deployment without retraining.
These results demonstrate that the proposed deep semantic iterative reconstruction enables a single universal model to achieve strong cross-domain generalization and robust normality modeling under both extreme data scarcity and full-data conditions.
@@ 4.7 Anomaly Map Visualization
To qualitatively evaluate the localization capability of DSIR, we visualize the anomaly maps generated under the one-shot universal setting. For each dataset, we randomly select representative normal and abnormal samples from the full test set. Following the exact inference pipeline, we compute four anomaly maps per sample: the fused map from Loop 1, Loop 2, Loop 3, and the final fused sum across all loops.
@@
Figure 2 presents typical visualization results. For normal samples, all four maps remain uniformly clean across loops, demonstrating that DSIR correctly reconstructs normal semantics. For abnormal samples, the anomaly regions become progressively sharper and more localized as the number of loops increases, with the final fused map providing the clearest boundary and highest contrast. These visualizations confirm that the multi-loop iterative refinement not only improves quantitative AUC but also yields semantically meaningful and human-interpretable anomaly localization.
@@ 5 Conclusion
In this paper, we presented Deep Semantic Iterative Reconstruction (DSIR), a framework for unsupervised medical anomaly detection. DSIR trains a single universal model on only one normal image from each of nine heterogeneous datasets, enabling direct anomaly detection across all corresponding test sets without task-specific retraining. The framework introduces a simple re-downsampling decoder that performs semantic reconstruction of hierarchical teacher features, combined with a multi-loop iterative refinement mechanism that progressively strengthens normality constraints in deep feature space.
Extensive experiments across nine diverse medical benchmarks demonstrate that DSIR achieves state-of-the-art image-level AUC under the extremely limited one-shot universal setting, consistently outperforming both full-shot and few-shot baselines. The method also maintains strong performance when more normal samples become available and delivers competitive results in conventional per-dataset scenarios. Ablation studies confirm that the iterative refinement and multi-scale fusion are essential for robust cross-domain generalization.
DSIR offers a practical and scalable solution for real-world clinical applications where annotated data and large normal sets are unavailable.
@@
Although DSIR achieves strong image-level detection through deep semantic reconstruction, its pixel-level localization remains limited by the inherent abstraction of feature-space modeling. Highly irregular lesion boundaries and subtle intensity variations may not be delineated with perfect precision. Future research could incorporate explicit pixel-level constraints, multi-scale attention mechanisms, or integration with foundation models to further refine localization accuracy while preserving the universal one-shot capability.